# Low-Cost Architecture and Efficient Pattern Synthesis for Polarimetric Phased Array Based on Polarization Coding Reconfigurable Elements

### v2

@@ Abstract
Polarimetric phased arrays (PPAs) enhance radar target detection and anti-jamming capabilities. However, the dual transmit/receive (T/R) channel requirement leads to high costs and system complexity. To address this, this paper introduces a polarization-coding reconfigurable phased array (PCRPA) and associated pattern synthesis techniques to reduce PPA costs while minimizing performance degradation. Each PCRPA element connects to a single T/R channel and incorporates two-level RF switches for real-time control of polarization states and waveforms. By adjusting element codes and excitation weights, the PCRPA can generate arbitrarily polarized and dual-polarized beams. Efficient beam pattern synthesis methods are also proposed, featuring novel optimization constraints derived from theoretical and analytical analysis of PCRPAs. Simulations demonstrate that the approach achieves low cross-polarization and sidelobe levels comparable to conventional architectures within the scan range, particularly for large arrays. However, the channel reduction inevitably incurs power and directivity loss. Experiments conducted on an $8\times 8$ X-band array antenna validate the effectiveness of the proposed system. The PCRPA and synthesis methods are well-suited for large-scale PPA systems, offering significant cost-effectiveness while maintaining good sidelobe suppression and polarization control performance.
Index Terms:
Polarimetric phased array radar, low-cost phased array, polarization pattern synthesis, cross-polarization level (XPL).
@@ I Introduction
Polarimetric phased arrays (PPA) tend to be adopted by many radar systems to enhance their measurement performance within complex environments. The extra dimension of polarization can provide detailed information to detect targets in the presence of strong clutter[1, 2, 3, 4, 5], interference[6, 7, 8, 9, 10], or other adjacent objects[11, 12, 13]. These applications rely on the beamforming capabilities of dual- or arbitrarily polarized patterns, which require independent channels of orthogonal polarization states, usually horizontal (H) and vertical (V). Therefore, conventional PPA architectures generally consist of dual-polarized antenna elements with transmit/receive (T/R) channels connected to each of their ports[14, 15, 16, 17]. Compared with single-polarized phased arrays of the same size, the number of channels in the T/R modules of PPAs is doubled.
@@
The conventional architectures are high-configuration but add to the burden of manufacturing the PPA systems, especially for large-scale arrays. A typical cost distribution for phased arrays given by [18] shows that the T/R modules can account for half the system cost, whose components grow proportionally with the number of elements. Meanwhile, many recently reported advanced PPA radar systems are on the scale of 1000-element or even bigger[19, 20, 21] to obtain high space resolution and more degrees of freedom (DOFs) for signal processing. Therefore, for these large-scale PPA systems, the heavy use of T/R channels under conventional architectures will make the total cost and complexity fairly high. And this also reveals the practical opportunity for cost reduction.
@@
Over the past few decades, techniques like sparse arrays, thinned arrays, and subarrays have been thoroughly studied. They can significantly reduce the cost and complexity of phased arrays and directly apply to PPAs[22, 23]. However, these array designs are relatively fixed and static, so the configuration must be determined before manufacturing and cannot be changed. This will limit some crucial performance of radar systems, especially the power and scan range. In addition, polarization reconfigurable antennas are generally designed with single ports, requiring half the T/R channels compared to dual-polarized antennas[24, 25, 26]. If they are employed as PPA elements, the system cost and complexity will also be significantly reduced. Novel algorithms considering element polarization as variables can be proposed to ensure the performance of polarization pattern synthesis during the working process. Since the polarization states are adjusted in real-time, the design can be regarded as a dynamic thinned array with polarization as a new DOF.
@@
Current research related to this issue can be divided into two schemes. The first has thoroughly investigated the application of polarization-reconfigurable elements in metasurfaces[27, 28], transmitarrays[29], and other non-electronically scanned array antennas[30, 31]. These works are instructive, but their array architectures and pattern synthesis models differ from the PPAs. The second scheme explores the idea of mixed polarized elements in PPA pattern synthesis but does not involve specific architecture design or beamforming requirements[32, 33, 34, 35, 36]. For example, in [33, 34], a signal processing method called cross-polarization canceller (XPC) uses a selection of cross-polarized elements to improve the cross-polarization level (XPL) of beam patterns by canceling the corresponding field components. However, it is based on the conventional architecture and requires the phased array to be fully digital. Several architecture models have been reported in [35, 36], but they do not cover the commonly used modes of PPA radars, such as alternately transmitting and simultaneously receiving (ATSR) and simultaneously transmitting and simultaneously receiving (STSR).
@@
This paper proposes an architecture and corresponding polarization pattern synthesis methods for PPA radars to reduce the cost and complexity while ensuring beamforming performance. The proposed architecture employs elements with polarization reconfigurable ability. Given that the polarization states of each element are limited and independently controllable, this architecture is termed the polarization coding reconfigurable phased array (PCRPA). In the hardware design, each T/R channel has two switches: the antenna-connected switch for the antenna polarization state selection and the transceiver-connected switch for the waveform selection. This paper proposes novel constraints, derived from antenna radiation and array geometry analysis, to optimize pattern performance and enhance algorithm efficiency. Combined with classical optimization algorithms, efficient synthesis methods for the polarization pattern of the PCRPA are proposed. The main contributions of this paper are listed below.
@@
    A novel array architecture considering multiple applications for PPA radars is proposed. The designed architecture enables the real-time coding of polarization states and waveforms for reconfigurable antenna elements, allowing for the synthesis of arbitrarily polarized and dual-polarized beam patterns. Compared to conventional architectures, the proposed design offers a cost-effective advantage by reducing the number of active channels in the T/R modules by half.
@@
The remainder of this paper is organized as follows. Section II introduces the general design of the PCRPA and its beamformer architecture. Section III proposes the methods of polarization pattern synthesis based on the PRCA and verifies them through simulations. Section IV analyzes the parameter sensitivity of pattern synthesis, including the parameters of polarization state, array size, and beam position. Section V presents the experiment results to verify further the effectiveness of the proposed architecture and synthesis methods concerning PPA radars, then discusses the applicable scenarios. Section VI concludes the entire paper.
@@ II General Architecture of The PCRPA
Figure 1: The illustration of the conventional and the proposed polarimetric phased array architectures. (a) Architecture of the conventional polarimetric phased array, where two polarization ports of each element are connected to independent channels. (b) Architecture of the proposed polarimetric phased array with polarization reconfigurable elements. (c) Schematic diagram of the proposed array, which can produce dual-polarized or arbitrarily polarized beams by controlling the polarization states and waveforms of elements.
The architecture designs of the conventional PPA and the proposed PCRPA are illustrated in Fig. 1a and 1b. Although both diagrams adopt uniform $N$-element linear arrays with the spacing $d$ for simplicity, the proposed architecture can also be utilized in various antennas, including planar and conformal arrays. The elements for the conventional PPA are dual-polarized antennas consisting of two orthogonal polarization ports (usually H and V polarization), represented by the red and blue rectangles in Fig. 1a. Regardless of the subarray design, each port is connected to an independent T/R channel, corresponding to the independent weighting $a_{ih}\phi_{ih}$ or $a_{iv}\phi_{iv}$ and waveform H or V. In this case, the electric field of beams formed by H and V ports can be expressed as:
@@
where $\boldsymbol{e}_{ih}(\theta,\phi)$ and $\boldsymbol{e}_{iv}(\theta,\phi)$ denote the electric field vectors of the $i$-th H and V ports, and the effects of amplitude weights $a_{ih}$, $a_{iv}$ and phase weights $\phi_{ih}$, $\phi_{iv}$ have already been included.
For the proposed PCRPA architecture, the elements are polarization reconfigurable and represented by the spliced rectangles in Fig. 1b. The colored and transparent halves of the rectangles indicate polarization excited and unexcited, respectively. These elements are each connected to a single T/R channel, corresponding to the independent weighting $a_{i}\phi_{i}$ and waveform H or V. Compared to the conventional PPA, the number of T/R channels here is reduced to half. The PCRPA can form multiple polarized beams by controlling the coding states of elements, where the beam can be expressed as:
@@
where $\boldsymbol{E}_{j}(\theta,\phi)$ is the electric field of the $j$-th beam, $\boldsymbol{I}^{jh}$ and $\boldsymbol{I}^{jv}$ are vectors containing the element indices coded by H and V polarization for the $j$-th beam, $N_{jh}$ and $N_{jv}$ are the number of these elements. Here, vectors $\boldsymbol{I}^{jh}$ and $\boldsymbol{I}^{jv}$ should satisfy the constraints followed:
$\left\{{\begin{array}[]{l}{1\leq I^{jA}_{i}\leq N,\>I^{jA}_{i}\in\mathbb{Z},i=1,2,\cdots,N_{jA}}\\ {I^{jA}_{m}\neq I^{jA}_{n},\>\forall m,n\in\{1,2,\cdots,N_{jA}\},m\neq n}\\ {\boldsymbol{I^{jh}\cap\boldsymbol{I}^{jv}=\emptyset}}\end{array}}\right$. (3)
@@
Therefore, it can be summarized from Fig. 1b and Equation (2)-(3) that beam patterns of the PCRPA are synthesized by 3 groups of variable: the amplitude and phase weights, the waveform code and the polarization code. The latter two groups exist only in the proposed architecture rather than the conventional PPA. By adjusting these variables in real-time, the PCRPA can form multiple electronically scanned beams with any desired polarization. Considering the practical scenarios in radar applications, this paper focuses on two problems: arbitrarily polarized and dual-polarized pattern synthesis, shown in Fig. 1c. These beam patterns of high performance can meet the requirements of polarization measurements and anti-interference.
Since the proposed PCRPA architecture includes a two-level coding process, it can be realized by the beamformer designed with two-level RF switches shown in Fig. 2. Here, RF switches are utilized to connect the array antenna, T/R modules, and transceivers sequentially. To highlight the architectural features, the radar transmitters and receivers are schematically represented in the diagram. Regarding digitization levels: The current design reflects the active phased array architecture; Subarray digital architecture can be achieved by incorporating multiple digital transceivers after the T/R modules; For fully digital architectures, the second-level switches become redundant and can be eliminated[37]. This is because the waveforms can be conveniently altered by direct digital synthesizers (DDS) in digital transceivers at the element level. Besides, the combination of dual-polarized antennas and RF switches can also be replaced by polarization reconfigurable antennas with single ports. However, the latter may not tolerate high transmitting power due to components like PIN diodes.
Figure 2: The beamformer design of the proposed phased array architecture for active arrays or digitized subarrays, where RF switches are utilized between T/R modules, array antennas, and transceivers.
@@ III-A Problem Formulation
The beam patterns are formed for the active phased arrays by adding each AEP with independent weights. In this case, the array pattern can be expressed as[38]:
@@ III-B Arbitrarily Polarized Pattern Synthesis
For arbitrarily polarized pattern synthesis problems, define the desired polarization basis according to Ludwig’s polarization definition II[39]:
@@
Based on Equation (13), the conventional PPAs can synthesize patterns of arbitrarily polarized beams by directly multiplying the corresponding weights $\cos{\gamma}$ and $\sin{\gamma}e^{-j\eta}$ into the polarization channels H and V[36]. However, this method ignores the cross-polarization components radiated at the H and V ports. If they are considered, there will be:
@@
The whole method can be summarized step by step, given in Algorithm 1. Although Equation (18) can uniquely calculate the number of H and V elements, this method provides an adjustment mechanism realized by Steps 4 and 7. This is because the AEPs of embedded elements may not fully represent the array polarization characteristics due to mutual coupling and other effects. For each $n_{h}$ and $n_{v}$, a maximum generation is set. If the iteration reaches the maximum generation but the patterns do not meet the requirements, the method will adjust $n_{h}$ and $n_{v}$. The adjustment rules can be flexible, for example, reducing one H element each time. The random number generator in Step 5 is chosen according to simulation results given by Fig. 4, which can also be replaced by other optimization algorithms if needed.

### v3

@@ Abstract
Polarimetric phased arrays (PPAs) enhance radar target detection and anti-jamming capabilities, but their conventional dual transmit/receive (T/R) channel architecture leads to high cost and system complexity. To address these limitations, this paper proposes a polarization-coding reconfigurable phased array (PCRPA) and associated pattern synthesis techniques, which reduce the channel count while preserving key performance. In the PCRPA, each antenna element connects to a single T/R channel and is equipped with a two-level RF switch, enabling real-time control of its polarization state and subarray grouping. By optimizing both the element polarization codes and the excitation weights, the array can synthesize arbitrarily polarized and dual-polarized beams. Simulation results show that the proposed approach achieves suppressed cross-polarization and comparable sidelobe levels compared to conventional PPAs across a wide scan range, with performance improvements being more pronounced in larger arrays. The inherent channel reduction does, however, incur a trade-off in terms of radiated power and directivity. Experimental validation using an $8\times 8$ X-band array antenna confirms the feasibility and effectiveness of the proposed system. The PCRPA architecture and the accompanying synthesis methods offer a cost-effective solution for large-scale PPA systems, maintaining sidelobe and polarization control with significantly reduced hardware complexity.
@@ I Introduction
Polarimetric phased arrays (PPA) tend to be adopted by many radar systems to enhance their measurement performance within complex environments. The extra dimension of polarization can provide detailed information to detect targets in the presence of strong clutter[hurtado2008polar, li2014ground, cheng2015antenna, wu2024a, an2024ground], interference[compton1981on, lu2021adaptive, yin2022radio, pan2023joint, an2024radio, xie2025coarray], or other adjacent objects[zhu2024principles, zhu2024a, liu2024robust]. These applications rely on the beamforming capabilities of dual- or arbitrarily polarized patterns, which require independent channels of orthogonal polarization states, usually horizontal (H) and vertical (V). Therefore, conventional PPA architectures generally consist of dual-polarized antenna elements with transmit/receive (T/R) channels connected to each of their ports[salazar2015transmit, ta2015crossed, salazar2015tr, mirmozafari2019dual]. Compared with single-polarized phased arrays of the same size, the number of channels in the T/R modules of PPAs is doubled.
@@
Conventional architectures offer high performance but significantly increase the manufacturing cost and complexity of PPA systems, especially for large-scale arrays. A typical cost distribution for phased arrays given by [herd2016the] shows that the T/R modules can account for half the system cost, whose components grow proportionally with the number of elements. Meanwhile, many recently reported advanced PPA radar systems are on the scale of 1000-element or even bigger[fulton2017cylindrical, zhang2022cylindrical, palmer2023horus] to obtain high space resolution and more degrees of freedom (DoFs) for signal processing. Therefore, for these large-scale PPA systems, the heavy use of T/R channels under conventional architectures will make the total cost and complexity fairly high, which also reveals the practical opportunity for cost reduction.
@@
Over the past few decades, techniques like sparse arrays, thinned arrays, and subarrays have been thoroughly studied. They can significantly reduce the cost and complexity of phased arrays and directly apply to PPAs[rocca2016uncon, mailloux2009irreg]. However, these array designs are relatively fixed and static, so the configuration must be determined before manufacturing and cannot be changed. This will limit some crucial performance of radar systems, especially the power and scan range. In addition, polarization reconfigurable antennas are generally designed with single ports, requiring half the T/R channels compared to dual-polarized antennas[nishamol2011an, lin2017multi, chen2021a]. If they are employed as PPA elements, the system cost and complexity will also be significantly reduced. Novel algorithms considering element polarization as variables can be proposed to ensure the performance of polarization pattern synthesis during the working process. Since the polarization states are adjusted in real-time, the design can be regarded as a dynamic thinned array with polarization as a new DoF.
@@
Current research related to this issue can be divided into two schemes. The first has thoroughly investigated the application of polarization-reconfigurable elements in metasurfaces[yang2021recon, yang2023a], transmitarrays[wan2023code], and other non-electronically scanned array antennas[chen2022, wang2022]. These works are instructive, but their array architectures and pattern synthesis models differ from the PPAs. The second scheme explores the idea of mixed polarized elements in PPA pattern synthesis but does not involve specific architecture design or beamforming requirements[barbetty2012, salazar2024, salazar2024a, wang2024low-cost, zhou2024]. For example, in [salazar2024, salazar2024a], a signal processing method called cross-polarization canceller (XPC) uses a selection of cross-polarized elements to improve the cross-polarization level (XPL) of beam patterns by canceling the corresponding field components. However, it is based on the conventional architecture and requires the phased array to be fully digital. Several architecture models have been reported in [wang2024low-cost, zhou2024], but they do not cover the commonly used modes of PPA radars, such as alternately transmitting and simultaneously receiving (ATSR) and simultaneously transmitting and simultaneously receiving (STSR).
@@
This paper proposes an architecture and corresponding polarization pattern synthesis methods for PPA radars to reduce the cost and complexity while ensuring beamforming performance. The proposed architecture employs elements with polarization reconfigurable ability. Given that the polarization states of each element are limited and independently controllable, this architecture is termed the polarization coding reconfigurable phased array (PCRPA). In the hardware design, each T/R channel has two switches: the antenna-connected switch for the antenna polarization state selection and the transceiver-connected switch for the subarray selection. This paper proposes novel constraints, derived from antenna radiation and array geometry analysis, to optimize pattern performance and enhance algorithm efficiency. Combined with classical optimization algorithms, efficient synthesis methods for the polarization pattern of the PCRPA are proposed. The main contributions of this paper are listed below.
@@
    A novel array architecture considering multiple applications for PPA radars is proposed. The designed architecture enables the real-time coding of polarization states and subarrays for reconfigurable antenna elements, allowing for the synthesis of arbitrarily polarized and dual-polarized beam patterns. Compared to conventional architectures, the proposed design offers a cost-effective advantage by reducing the number of active channels in the T/R modules by half.
@@
The remainder of this paper is organized as follows. Section II details the architecture design of the proposed PCRPA. Section III presents the theoretical foundation, formulating the efficient pattern synthesis methods. Section IV provides a comprehensive simulation validation and performance analysis, including parameter sensitivity studies. Section V offers experimental verification on a physical array prototype, discussing the agreement between measured and simulated results. Finally, Section VI concludes the paper.
@@ II General Architecture of The PCRPA
Figure 1: Illustration of the conventional and the proposed PPA architectures. (a) Conventional architecture. (b) Proposed architecture. (c) Schematic diagram of the proposed architecture, which can produce dual-polarized or arbitrarily polarized beams by controlling the polarization states and subarrays of elements.
This section begins with an overview of the conventional PPA architecture. As illustrated in Fig. 1a, the architecture consists of $N$ elements spaced at a uniform distance $d$ on the YOZ plane. In this architecture, the H and V polarization ports of each dual-polarized antenna element are connected to independent T/R channels (demonstrated by weight $a_{ih}\phi_{ih}$ or $a_{iv}\phi_{iv}$) and are fixedly assigned to corresponding H-polarized or V-polarized subarrays. The system synthesizes beams independently within the H and V subarrays via separate beamforming networks, which are then coherently combined in the far-field to achieve dual-polarized beam calibration or to synthesize beams with arbitrary polarization states. In this case, the electric field of beams formed by H and V subarrays can be expressed as:
@@
where $\boldsymbol{e}_{ih}(\theta,\phi)$ and $\boldsymbol{e}_{iv}(\theta,\phi)$ denote the electric field vectors of the $i$-th H and V ports, and the effects of amplitude weights $a_{ih}$, $a_{iv}$ and phase weights $\phi_{ih}$, $\phi_{iv}$ have already been included. However, this conventional approach requires a large number of TR channels, leading to high system cost and complexity.
To address this issue, this paper proposes a PCRPA architecture, as illustrated in Fig. 1b, where the number and positions of elements remain consistent with those in Fig. 1a. The idea for cost reduction is achieved through the resource sharing of T/R channels among the polarization ports and subarray modules. Specifically, polarization-reconfigurable antenna elements are employed, allowing the dual-polarized ports of each element to time-share a single T/R channel, thereby halving the total number of channels. Furthermore, to address the beam reconfiguration requirements for dual-polarized pattern synthesis, a similar approach is applied, enabling the H- and V-polarized subarrays to also time-share a T/R channel. These two levels of reconfiguration collectively form an array architecture in which both the polarization and subarray connections are dynamically reconfigurable.
In this way, the PCRPA architecture realizes beam synthesis through a dual-coding mechanism. First, subarray coding determines the set of element indices, denoted as $\boldsymbol{I}_{j}$, excited for the $j$-th beam. Second, polarization coding defines the sets of indices for all elements in the array excited through their H ports and V ports, denoted as $\boldsymbol{I}_{h}$ and $\boldsymbol{I}_{v}$. The total radiation pattern of beam $j$ is the superposition of the fields radiated by the elements in the intersection set of $\boldsymbol{I}_{jh}$ and $\boldsymbol{I}_{jv}$, defined as:
$\left\{{\begin{array}[]{l}{\boldsymbol{I}_{jh}=\boldsymbol{I}_{j}\cap\boldsymbol{I}_{h}}\\ {\boldsymbol{I}_{jv}=\boldsymbol{I}_{j}\cap\boldsymbol{I}_{v}}\end{array}}\right$. (2)
And the electric field of the $j$-th beam can be expressed as:
@@
where $N_{jh}$ and $N_{jv}$ are the number of elements in sets $\boldsymbol{I}_{jh}$ and $\boldsymbol{I}_{jv}$. Here, vectors $\boldsymbol{I}^{jh}$ and $\boldsymbol{I}^{jv}$ should satisfy the constraints followed:
$\left\{{\begin{array}[]{l}{1\leq I^{jA}_{i}\leq N,\>I^{jA}_{i}\in\mathbb{Z},i=1,2,\cdots,N_{jA}}\\ {I^{jA}_{m}\neq I^{jA}_{n},\>\forall m,n\in\{1,2,\cdots,N_{jA}\},m\neq n}\\ {\boldsymbol{I}^{jh}\cap\boldsymbol{I}^{jv}=\varnothing}\end{array}}\right$. (4)
@@
Based on Equation 3, the proposed PCRPA architecture is able to form electrially scanned beams of either arbitrary or dual polarization, shown in Fig. 1c. This is owing to two key aspects: Firstly, the code vectors $\boldsymbol{I}_{jh}$ and $\boldsymbol{I}_{jv}$ create irregular distributions of phase centers across the aperture, analogous to the synthesis principles of sparse arrays, which enables the formation of high-gain, low-sidelobe beams. Secondly, by dynamically configuring the number and excitation coefficients (amplitude and phase) of elements fed through different polarization ports, the architecture can achieve functionality similar to cross-polarization calibration techniques, allowing for the precise synthesis of any desired polarization state.
Figure 2: The beamformer design of the proposed phased array architecture for different phased array architectures. (a) For subarray-level digital phased array. (b) For fully digital phased array.
TABLE II: Comparison of PPA Architectures
Architecture Channel Number Element Polarization Beamforming DoFs Arbitrarily Polarized Beamforming Dual Polarized Beamforming
Conventional $2N$ Fixed Weight Yes Yes
Ref. [zhou2024] NN Fixed Weight Yes No
Proposed NN Reconfigurable Weight, Port, Subarray Yes Yes
In terms of the hardware implementation, the two-level coding process of the PCRPA can be realized by the two-level of RF switches shown in Fig. 2a. Here, RF switches are utilized to connect the array antenna, T/R modules, and beamformers sequentially. Here, the combination of dual-polarized antennas and RF switches may also be replaced by polarization reconfigurable antennas with single ports. However, the latter may not tolerate high transmitting power of radar systems at presence due to components like PIN diodes. Besides, although this paper focuses on the subarray-level digital systems, the PCRPA is also applicable to fully digital phased arrays. In a fully digital architecture shown by Fig 2b, the subarray coding function can be implemented directly by the digital beamformer[fulton2016digital], allowing the second-level RF switch to be omitted, thereby simplifying the hardware structure.
To summarize, a comparative analysis is provided by Table II. In comparison to the conventional PPAs, both the proposed PCRPA architecture and a previously reported structure by the team [zhou2024] can reduce the number of T/R channels by half, lowering the system cost and complexity. Relative to the prior work, the key difference of the PCRPA lies in beamforming DoFs and capabilities. While [zhou2024] utilized fixed phase shifters to decide element polarization, this work adopted a real-time reconfigurable design. This makes the port polarization a new DoF for beamforming in the proposed architecture. Additionally, the subarray coding DoF is introduced. These enhancements enable the PCRPA to achieve dual-polarized beamforming. Meanwhile, the real-time reconfigurability of polarization coding also leads to better radiation performance, which will be demonstrated later.
@@ III-A Problem Formulation
The beam patterns are formed for the active phased arrays by adding each AEP with independent weights. In this case, the array pattern can be expressed as[pozar1994the]:
@@ III-B Arbitrarily Polarized Pattern Synthesis
For arbitrarily polarized pattern synthesis problems, define the desired polarization basis according to Ludwig’s polarization definition II[ludwig1973the]:
@@
Based on Equation (14), the conventional PPAs can synthesize patterns of arbitrarily polarized beams by directly multiplying the corresponding weights $\cos{\gamma}$ and $\sin{\gamma}e^{-j\eta}$ into the polarization channels H and V[zhou2024]. However, this method ignores the cross-polarization components radiated at the H and V ports. If they are considered, there will be:
@@
The proposed algorithm, outlined in Algorithm 1, takes the AEPs of all ports, along with performance specifications (e.g., beam steering, polarization, and sidelobe level), and outputs the polarization code and weight vectors accroding to Fig. 3. A central challenge addressed by this method is the mitigation of mutual coupling effects between elements and ports. To this end, the algorithm incorporates a two-fold strategy: First, while the polarization basis in Step 2 is constructed using the data of the center embedded element, the array pattern synthesis in Step 5 is computed based on the full AEP model. This model utilizes the full-wave simulated or measured data of all ports, thereby accounting for mutual coupling. Second, to compensate for any residual inaccuracies arising from the single-element estimation in Step 2, Step 7 provides an adjustment mechanism. Specifically, if the pattern requirements are not met upon reaching the maximum generation for a given $n_{h}$ and $n_{v}$, the ratio of H-to-V elements is adaptively modified according to a predefined rule (e.g., reducing one H element per adjustment). Furthermore, the random number generator in Step 5, whose distribution is informed by the Monte Carlo simulation results in Fig. 4, can be seamlessly replaced by other optimization algorithms for initial value generation if required.